OpenStudy (anonymous):

how to get dy/dx for (2^-y) + (2^-x)=1

OpenStudy (anonymous):

HINT: Logarithm

OpenStudy (anonymous):

is it can ther was a "+' ???

OpenStudy (anonymous):

is the answer -1??

OpenStudy (anonymous):

Wait, we want to derivative y, so first we have to isolate y. Can you do it?

OpenStudy (anonymous):

It's going to be ugly, trust me.

OpenStudy (anonymous):

mmm i hv a question .. i think only include ln when "*" or "/" is it incorrect??

OpenStudy (anonymous):

Yeah. ln(a·b) = ln a + ln b ln(a/b) = ln a - ln b I forgot about that, that's why I told you to do something else.

OpenStudy (anonymous):

is it incorrect?

OpenStudy (anonymous):

No, you are correct.

OpenStudy (anonymous):

so how inlude ln here?

OpenStudy (anonymous):

Forgot ln, it's probably won't help. To solve your problem, either derivative both sides then isolate y' or isolate y then derivative it. Can you do either of them?

OpenStudy (anonymous):

can u tell me a way ?

OpenStudy (anonymous):

Either derivative both sides then isolate y' or isolate y then derivative it. Can you do one of them?

OpenStudy (anonymous):

so -yln(2)-xln(2) =0 y=-x dy/dx= -1 is it correct?

OpenStudy (anonymous):

incorrect??

OpenStudy (anonymous):

No, there is no numerical answer. We have two unknown variables. We just need to find y' or dy/dx

OpenStudy (anonymous):

2^-y +2^-x =1 ln{2^-y +2^-x }=ln 1 -2^y ln(2)-2^xln(20=0 2^y=-2^x y=-x dy/dx=-1 ???????

hartnn (hartnn):

derivative of a^x is ?

hartnn (hartnn):

u need to know that formula and CHAIN rule,.... u know chain rule ??

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

what @hartnn said you need both \(\frac{d}{dx}b^x=b^x\ln(b)\) and the chain rule to get \(\frac{d}{dx}2^{-x}=-2^{-x}\ln(2)\)

OpenStudy (anonymous):

can u explain it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the derivative of \(b^x\) is \(b^x\ln(b)\) so for example the derivative of \(10^x\) is \(10^x\ln(10)\) i can explain why this is true if you like

OpenStudy (anonymous):

is it {2^(x+y) -2^x} / {2^y - 2^ (x+y) }

OpenStudy (anonymous):

is it the answer?

OpenStudy (anonymous):

to know why this is true, you need to know that \(b^x=e^{x\ln(b)}\) and then you can take the derivative using the chain rule, you get \(e^{x\ln(b)}\times \ln(b)\) which is the same as \(b^x\ln(b)\)

OpenStudy (anonymous):

i doubt it, but we can check start with \[ 2^{-y} + 2^{-x}=1\] take derivative wrt \(x\) and get \[-\ln(2)2^{-y}y'-\ln(2)2^{-x}=0\]

OpenStudy (anonymous):

now solve for \(y'\) you do not take the log of the whole thing which i think is confusing you you cannot "simplify" \(2^{-y}+2^{-x}\) by taking the log . it is a sum, not a product

OpenStudy (anonymous):

how u include ln ? there was "+"????

OpenStudy (anonymous):

i think there are a few things confusing here lets start with the basics do you know what the derivative of \(2^{-x}\) is? you need that one

OpenStudy (anonymous):

2^x is 2^x ln(2) 2^y is 2^y dy/dx ln(20 2^(x+y) is 2^(x+y) (1+dy/dx) ln(2)

OpenStudy (anonymous):

ok but you do not have \(2^{x+y}\) in the problem

OpenStudy (anonymous):

so 2^x + 2^y dy/dx = 2^(x+y) (1+dy/dx)

OpenStudy (anonymous):

the original problem is this; 2^x +2^y = 2^(x+y)

OpenStudy (anonymous):

i thought you had \[2^{-y}+2^{-x}=1\] ??

OpenStudy (anonymous):

its correct too it was divide by 2^(x+y)

OpenStudy (anonymous):

ok then lets start with what you wrote above do you know the derivative of \(2^{-x}\)?

OpenStudy (anonymous):

it is -x* 2^-x * ln(2)

OpenStudy (anonymous):

no it is \(-\ln(2)2^{-x}\)

OpenStudy (anonymous):

oh yes not x sorry

OpenStudy (anonymous):

so -ln2^-2 -dy/dx ln(2)2^-y =0

OpenStudy (anonymous):

so then we are ready to go you get \[-\ln(2)2^{-y}\times y'-\ln(2)2^{-x}=0\]solve for \(y'\)

OpenStudy (anonymous):

dy/dx= 2^-x / 2^-y ?????

OpenStudy (anonymous):

off by a minus sign

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

but the given answer is; {2^(x+y) - 2^x} / {2^y - 2^(x+y)}

OpenStudy (sirm3d):

how about this: \[2^{-y}+2^{-x}=1\] or \[2^x+2^y=2^{x+y}\] after removing negative exponents. now apply the derivative to both sides.